OpenStudy (anonymous):

if the second term of an arithmetic sequence is 10 and the seventh term is 35 find the 11th term?

OpenStudy (anonymous):

let the first term be a and the common difference be d.then for second term 10=a+d and for the seventh term 35=a+6d .now u have two unknown and two equations.u can solve it easily